OpenStudy (anonymous):

What is the sign of the product (3)(−25)(7)(−24)?

OpenStudy (yanasidlinskiy):

\(\Huge\bf \color{green}{Welcome~to~OpenStudy!!}\hspace{-310pt}\color{cyan}{Welcome~to~OpenStudy!!}\hspace{-307.1pt}\color{green}{Welcome~to~\color{green}{Open}}\color{green}{Study!!!!}\) Hint:!!! Product = what?

OpenStudy (uri):

Use the rule of PEMDAS. @jolorido

OpenStudy (anonymous):

im still confused lol

OpenStudy (anonymous):

Sign of the product is positive: +3(-25)= -75 -75(+7)= -525 -525(-24)=+12,600 RULES: positive times positive = positive positive times negative = negative

OpenStudy (anonymous):

negative times negative = positive

OpenStudy (anonymous):

@YanaSidlinskiy So you say "Welcome to OpenStudy" then leave the question without answering it? All I can say is that is rude.

OpenStudy (anonymous):

PEMDAS is 4 steps, Parenthesis 1st. Exponents second, Multiplication and Division 3rd, (which ever comes first) and addition and subtraction (which ever comes first)